OpenStudy (anonymous):

Evaluate the expression if x=-3 -13x -39 16 39 10

OpenStudy (shamim):

x=-3-13x x+13x=-3 14x=-3 x=-3/14

OpenStudy (anonymous):

how come you got -13?

OpenStudy (anonymous):

anwer this: 3x-7